OpenStudy (anonymous):

can somebody show me the process of getting the slope intercept form and point slope form of A line passing through (4, -7) with a slope of ½?

OpenStudy (anonymous):

fot Point (x1,y1) and slope m we have (y-y1)=m(x-x1)

OpenStudy (anonymous):

for this data we get y=x/2 -9 which is slope intercept form and (y+7)=1/2(x-4) point slope form

OpenStudy (anonymous):

okay thanks!